OpenStudy (anonymous):

Identify the vertex and the axis of symmetry of the graph of the function y = 3(x + 2)^2 – 3. vertex: (2, –3); axis of symmetry: x = 2 vertex: (–2, –3); axis of symmetry: x = –2 vertex: (2, 3); axis of symmetry: x = 2 vertex: (–2, 3); axis of symmetry: x = –2

OpenStudy (anonymous):

a(x-h)^2+k does that help

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

y = 3(x + 2)^2 – 3. y = a(x - h)^2 + k

OpenStudy (anonymous):

vertex is (h,k)!

OpenStudy (anonymous):

thanks ! i get it